What is the time constant for the discharge of the capacitor in thefigure?
Answer #1
the equivalent resistance :
R = 2*1x10^3 = 2000 Ω
=> the time constant : T = R*C = 2000*1x10^-6 =2x10^-3 s = 2 ms
Answer #2
Given
     R1 = 1 kΩ
      R2 = 1 kΩ
C = 1 μF
R1 and R2 are connected in seriescombination  
          R = R1+ R2     =   2 kΩ  =200 Ω
       timeconstant     τ = R C    = 200 * 10-6    = 2*10-3   s
Answer #3
the time constant for the discharge through resistance of thecapacitor is:
T=RC
because euqation discharge



=> T=(2000*10^-6)=2*10^-3s